UT 20150229 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, WSeifert, APramskiy, DThompson (LBTO)
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator: Sallenson
LBTO Support: TSargent

Plan for the night

- If AO software workaround is successfull, do some AO-tests during first half
than do some SL-tests with LUCI2 during second half of night

Summary

  • Problem with AO/AGw electronics not solved

  • SW Patch as workaround implemented, loop can be closed (but is not very stable)

  • same AO-tests, double star with 0.35" separation easily resolved

Preparations

- Problem with AO/AGw electronics is due to electropiezo. Hardware issue which can not be solved immediately.
Simone et al. try to find a workaround by modifying their software

Details

All times are UT
Files are luci2.20150329.NNNN.fits

01:42 Sunset, sky mostly clear, T=7deg, Hum=14%, low wind

Simone et al. are still working on their SW modification

5:00 Some cirrus around, somewhat thicker clouds down in the west

5:50 AO SW modification done for now, Steve opens dome, checks pointing
and collimates on A45_R8_0035 (HD 104424, R+I ~ 8.2mag)

6:35 Trying to do an AO preset, need to reacquire star, redo AO preset, fixed ROTangle=210deg

Sky coverage by clouds increases => no AO correction on this star possible (4mag absorption)

07:10 Moving to 23 Com (R+I ~ 5mag), Steve needs to home the hexapod

07:37 AO preset on 23 Com, success, loop closed, gnagna...thick clouds coming in
slightly better, loop can not permanently be closed due to clouds. Star is double with 0.35" separation,
easily visible (see Fig.)!!! Input seeing 1.2". Strehl can not be computed since double star (background).

LUCI2_N30_FeII_23Com.jpg Double-300ms_NCPA_On.jpg

08:50 Preset to A45_R8_0037 (HD 112515), R+I ~ 8.1mag, loop closed

09:30 Sky mostly clear now

09:40 GEIRS frozen, can't (yet) determine Strehl, GEIRS fixed, no Strehl, A45_R8_0037 at too low elevation now

10:25 Preset to A45_R8_0045 (HD 146586), R+I ~ 8.1, collimation takes a while but didn't complete,
AO preset, loop closed (but again not very stable), seeing ~ 1.2"-1.4" now, Strehl about 50%

Sky completely clear, setting up system for some SL-obs

12:05 ZP_BS9166_N3.75.xml, ZPs zJHKKs, #253works in principle, except that images are not loaded in ALADIN

12:40 ZP_BS9166_N30.xml, ZPs zJHKKs, doesn't work, "error to update detector focus" for z-filter,
changed to z-filter manually, this worked after third attempt

Moving telescope to zenith for some flats

12:57 J: 2.75sec, #254-263, ~ 4-6k, #264-73, 7-22k

13:02 Too bright for H-filter,K,Ks-filter

13:05 H2, 2.75sec, #274-283, 3k

13:07 Sunrise

13:08 Br_gam, 2.75sec, #284-293, 3k

13:11 H2, 2.75sec, #284-293, 2k

13:11 H2, 10sec, #294-303, 4k

13:13 H2, 10sec, #304-313, 3k

13:18 H2, 5sec, #314-323, 6k

13:20 H2, 5sec, #334-333, 10k

13:23 Br_gam, 5sec, #334-343, 6k

13:26 Br_gam, 5sec, #344-353, 7-k

NB:

13:30 Closing

-- %USERSIG{JochenHeidt - 2015-03-29}%
I Attachment Action Size Date Who Comment
Double-300ms_NCPA_On.jpgjpg Double-300ms_NCPA_On.jpg manage 83 K 29 Mar 2015 - 09:08 JochenHeidt Double star with separation of 0.35" (log-scale)
LUCI2_N30_FeII_23Com.jpgjpg LUCI2_N30_FeII_23Com.jpg manage 50 K 29 Mar 2015 - 09:03 JochenHeidt Double star with separation of 0.35"
Topic revision: r9 - 03 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback